OpenStudy (anonymous):

-3x-9 all divided by -3 Simplify choices x + 3 x − 9 −3x + 3 x − 3

OpenStudy (anonymous):

-3x/-3 = 1x and -9/-3 =3 since a negatove / by a negatove is a positve so the answer is x+3

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

8x-10 all divided by -2 Simplify

OpenStudy (anonymous):

 (-)(-)=+ (+)(+)=+ (+)(-)=- (-)/(-)=+ (+)/(+)=+ (+)/(-)=-